# Exploiting Global Graph Homophily for Generalized Defense in Graph Neural Networks

### v1

@@ 3.3.1 The Number of Paths
We first consider how to obtain the number of paths $n$. Taking the $k$th power of the adjacency matrix can provide the number of $k$-hop paths between node pairs, which can be represented as $\mathbf{A}^{k}_{ij}$. Therefore, we can model the $k$th-order global homophily index $h_{k}$ as follows:
@@ 3.4 Derivation and Analysis
To obtain the result of the aforementioned formula, we further derived and optimized it.
@@
Suppose edge $e_{kl}$ is removed. Then, within the matrix $\Delta\mathbf{A}$, only $\Delta A_{kl}=\Delta A_{lk}=-1$ while all other elements remain 0. Given this alteration, we can deduce the values at each position in the matrix $M^{\prime}\Delta M$, with the calculation process outlined as follows:
@@ 3.5.3 Matrix Computation Strategy
$\mathcal{J}_{ij}=\frac{X_{i}X_{j}^{T}}{1-\left(1-X_{i}\right)\left(1-X_{j}\right)^{T}}=\frac{X_{i}X_{j}^{T}}{1-\bar{X}_{i}\bar{X}_{j}^{T}}=\frac{X_{i}X_{j}^{T}}{\overline{\bar{X}_{i}\bar{X}_{j}^{T}}}$ (11)
@@ 4.1.1 Datasets
In this study, we primarily utilized three widely recognized datasets: Cora, Amazon Photo, and Amazon Computers. The selection of these three datasets was based on three main considerations:
@@ 4.2.1 Accuracy Evaluation
We initially assess the accuracy of node classification for various defense methods against non-target adversarial attacks. The adversarial attack dataset, generated in a prior section using Metattack and PGD, was employed for our experiments. Tables 2 present the mean accuracy along with the standard deviation under the two respective attack methods. Performance rankings are indicated in the upper right corner, with the top-performing results highlighted in bold. As observed from the tables, our approach consistently leads under diverse disturbance rates, demonstrating significant efficacy under higher levels of disturbance. Despite the decline in classification accuracy across all defense methods as the attack intensity increased, it is worth noting that Talos tended to maintain a high classification accuracy and showed strong robustness under higher attack intensity.

### v2

# Talos: A More Effective and Efficient Adversarial Defense for GNN Models Based on the Global Homophily of Graphs

@@ 3.3.1 The Number of Paths
We first consider how to obtain the number of paths $n$. Taking the $k$th power of the adjacency matrix can provide the number of $k$-hop paths between node pairs, which can be represented as $\mathbf{A}^{k}_{ij}$. Therefore, we can model the $k$th-order global homophily index $\mathrm{Hom}^{(k)}$ as follows:
@@ 3.4 Derivation and Analysis
Suppose the edge $e_{kl}$ is removed, resulting in a new adjacency matrix $\mathbf{A}^{\prime}$. Here, $\Delta\mathbf{A}=\mathbf{A}^{\prime}-\mathbf{A}$. For $\Delta\mathbf{A}$, $\Delta\mathbf{A}_{kl}=\Delta\mathbf{A}_{lk}=-1$, with all other values being 0. We further derived and optimized this result to achieve the final outcome.
@@
Where $M^{\prime}=(I-\alpha\mathbf{A}^{\prime})^{-1}$. Given this modification, we can determine the values at each position in the matrix $M^{\prime}\Delta\mathbf{A}M$. The calculation process is detailed as follows:
@@ 3.5.3 Matrix Computation Strategy
$\mathcal{J}_{ij}=\frac{\mathbf{x}_{i}\mathbf{x}_{j}^{T}}{1-\left(1-\mathbf{x}_{i}\right)\left(1-\mathbf{x}_{j}\right)^{T}}=\frac{\mathbf{x}_{i}\mathbf{x}_{j}^{T}}{1-\bar{\mathbf{x}}_{i}\bar{\mathbf{x}}_{j}^{T}}=\frac{\mathbf{x}_{i}\mathbf{x}_{j}^{T}}{\overline{\bar{\mathbf{x}}_{i}\bar{\mathbf{x}}_{j}^{T}}}$ (11)
@@ 4.1.1 Datasets
In this study, we primarily utilized three widely recognized datasets: Cora[21], Amazon Photo[22], and Amazon Computers. The selection of these three datasets was based on three main considerations:
@@ 4.2.1 Accuracy Evaluation
We initially assess the accuracy of node classification for various defense methods against non-target adversarial attacks. The adversarial attack dataset, generated in a prior section using Metattack and PGD, was employed for our experiments. Tables 2 present the mean accuracy along with the standard deviation under the two respective attack methods. Performance rankings are indicated in the upper right corner, with the top-performing results highlighted in bold. As observed from the tables, our approach consistently leads under diverse disturbance rates, demonstrating significant efficacy under higher levels of disturbance. Despite the decmidrule in classification accuracy across all defense methods as the attack intensity increased, it is worth noting that Talos tended to maintain a high classification accuracy and showed strong robustness under higher attack intensity.
@@ 5 Conclusion
{ack}
This project is supported by the National Natural Science Foundation of China under grant No. 62306328 and the science and technology innovation Program of Hunan Province under grant No. 2023RC3021.
Appendix A Hyperparameter Analysis
In this study, Talos has two hyperparameters: the number of edges removed ($n_{r}$) and the rate of information loss ($\alpha$).
A.1 Number of Edges Removed ($n_{r}$)
The choice of the number of edges removed is a critical hyperparameter. Removing too few edges may lead to poor defense performance, while removing too many may cause over-purification issues. Recent studies [13] indicate that GNNs rely on the graph’s mainstream structure, whereas adversarial perturbations are non-mainstream. Therefore, in clean graphs, slightly removing weakly homophilic edges shows little negative impact. However, if Talos exhausts heterophilic edges and continues removing homophilic parts, over-purification may occur. We addressed this issue using a validation set. Over-purification increases validation error by disrupting the mainstream structure. This signal helps us halt the purification process at a reasonable stage.
A.2 Rate of Information Loss ($\alpha$)
For $\alpha$, according to Equation 4, it should be within the range $[0,\frac{1}{\rho(\mathbf{A})}]$ for Talos to converge. Within this range, Talos performs stably, except near 0. Therefore, selecting a value slightly below $\frac{1}{\rho(\mathbf{A})}$ yields good results.
Regarding the calculation of $\rho(\mathbf{A})$, we know that computing $\rho(\mathbf{A})$ is resource-intensive for large graphs. Guo et al. [9] show that the spectral radius of a graph satisfies the following formula:
$\rho(\mathbf{A})\leq\sqrt{2|\mathcal{E}|-|\mathcal{V}|-k+\frac{5}{2}+\sqrt{2|\mathcal{E}|-2|\mathcal{V}|+\frac{9}{4}}}$
where $3\leq k\leq n$ is an integer that satisfies the following condition:
$\frac{(k-2)(k-3)}{2}\leq m-n\leq\frac{k(k-3)}{2}$
For small graphs, it is feasible to calculate the exact value of $\rho(\mathbf{A})$. However, for larger graphs, we can use the above formula to estimate $\rho(\mathbf{A})$ and thus select the hyperparameters accordingly.
Appendix B Comparing Mid-GCN and Talos Defense
DataSet Defense Method Attack Clean Graph Ptb5% Ptb10% Ptb15% Ptb20% Ptb25% Time
Cora Mid-GCN Metattack 82.31$\pm$0.47 79.80$\pm$0.61 76.03$\pm$0.71 69.51$\pm$1.09 60.48$\pm$1.51 59.11$\pm$1.86 0.73$\pm$0.10
PGD 82.28$\pm$0.53 78.88$\pm$0.49 76.29$\pm$0.42 74.27$\pm$0.57 72.79$\pm$0.66 71.37$\pm$0.56
Talos Metattack 82.71$\pm$0.49 80.83$\pm$0.74 78.31$\pm$0.76 74.17$\pm$0.39 68.19$\pm$0.98 65.28$\pm$1.16 0.60$\pm$0.10
PGD 81.06$\pm$0.91 78.81$\pm$0.47 78.42$\pm$0.54 75.97$\pm$0.41 75.40$\pm$0.64 73.98$\pm$0.34
Photo Mid-GCN Metattack 92.05$\pm$0.23 88.26$\pm$0.49 85.87$\pm$0.64 76.35$\pm$3.55 72.97$\pm$2.88 71.19$\pm$2.86 4.15$\pm$0.22
PGD 91.92$\pm$0.27 84.45$\pm$0.59 81.80$\pm$0.57 79.64$\pm$0.89 78.05$\pm$1.50 76.79$\pm$1.75
Talos Metattack 93.60$\pm$0.27 92.17$\pm$0.38 90.34$\pm$0.49 88.60$\pm$0.70 87.29$\pm$0.66 87.01$\pm$0.98 12.83$\pm$2.75
PGD 93.50$\pm$0.33 88.86$\pm$0.24 86.28$\pm$0.55 83.91$\pm$0.62 81.34$\pm$0.83 2.80$\pm$0.20
Computers Mid-GCN Metattack 85.62$\pm$0.34 80.02$\pm$0.47 73.47$\pm$0.89 72.19$\pm$0.93 70.67$\pm$0.96 69.81$\pm$1.37 12.83$\pm$2.75
PGD 85.64$\pm$0.32 77.25$\pm$0.85 72.40$\pm$0.71 69.33$\pm$0.83 67.58$\pm$0.82 65.75$\pm$1.19
Talos Metattack 88.98$\pm$0.39 85.58$\pm$0.53 78.30$\pm$1.91 80.58$\pm$1.05 80.49$\pm$0.67 79.83$\pm$1.73 12.83$\pm$2.75
PGD 88.98$\pm$0.39 85.29$\pm$0.36 82.95$\pm$0.69 81.03$\pm$0.91 79.07$\pm$0.86 11.90$\pm$1.54
Table B1: Performance Comparison of Mid-GCN and Talos under Adversarial Attacks
In recent research advancements, the Mid-GCN[11] model has been demonstrated as an effective method for countering adversarial attacks on graph neural networks (GNNs). Traditional defense strategies often rely on restoring the original graph structure or imposing robustness constraints during training to resist attacks. However, Mid-GCN employs an innovative approach by leveraging mid-frequency signals to enhance the model’s resilience against adversarial manipulations. Specifically, Mid-GCN integrates band-pass filters, allowing it to selectively emphasize graph signals that are less susceptible to noise and adversarial distortions—signals that are often overlooked by other methods.
To further evaluate the efficacy of Mid-GCN, we conducted comparative experiments with Talos, assessing their performance against various attacks, including Metattack and PGD, across multiple datasets such as Cora, Photo, and Computers. The experimental results, presented in Table B1, indicate that while Mid-GCN performs well in certain settings, Talos demonstrates superior protective capabilities in most high-disturbance scenarios. Particularly in high-perturbation conditions, Talos not only provides a more robust defense mechanism but also exhibits faster response times. This suggests that in complex attack environments, Talos is more effective and swift in safeguarding graph structures from adversarial influences, thereby maintaining the stability and accuracy of the model.
Appendix C Limitations
Despite the excellent performance of Talos in utilizing global homophily for adversarial defense in GNNs, its application scope has certain limitations. Specifically, Talos may not be effective when dealing with graphs exhibiting weak homophily. This limitation arises because the defense mechanism of Talos relies on the global homophily of the graph, which is insufficient to distinguish between normal and adversarial edges in graphs with weak homophily. To avoid the inappropriate use of Talos, we recommend measuring the homophily of a graph before applying Talos to ensure the effectiveness of the defense mechanism.